# From Embeddings to Dyson Series: Transformer Mechanics as Non-Hermitian Operator Theory

### 11 Mar 2026 v1

@@ Abstract
Transformer architectures are typically described in algorithmic and statistical terms, leaving their internal mechanics without a familiar structural language for researchers trained in physical theories. To bridge this gap, we develop a complementary operator-theoretic framework that recasts their mechanics in a language familiar to many-body physics. Beginning from the token as a discrete index without intrinsic geometry, we show that embedding corresponds to a basis transformation into a continuous representation space. Once such a reference basis is established, self-attention naturally assumes the role of a non-Hermitian interaction operator, and network depth implements an ordered composition of these interactions. Within this formulation, several empirical properties of deep Transformers—including stability at large depth, representational saturation, and the effectiveness of multi-head decomposition—find natural structural interpretations as consequences of regulated operator composition. Together, spectral geometry, channel factorization, and normalization emerge as organizing structural logic rather than isolated architectural choices. This perspective does not rely on post-hoc analogy, but follows a constructive path in which each parallel arises from the preceding structural step. By recasting Transformer mechanics in operator language, the framework lowers the conceptual barrier between deep learning and many-body physics through shared mathematical structure, making tools and intuitions from each domain more readily legible to the other.
@@ I Introduction: Two Different Universes
Modern many-body physics and deep learning originate from fundamentally different intellectual starting points. In quantum mechanics, a system is bounded by its Hamiltonian $H$ constructed from explicit dynamical laws, which defines its allowable states. Machine-learning systems, in contrast, are discovered through optimization: their structure emerges from data, with parameters adjusted to minimize a global loss over a statistical corpus whose generative rules are not explicitly encoded.
This difference in origin is not superficial. Physical operators inherit constraints, such as Hermiticity, unitarity, and conservation laws, which are non-negotiable consequences of physical law. On the other hand, learned operators inherit no such constraints as properties like stability are enforced through deliberate architectural choices rather than first principles. The degrees of freedom available to each framework are therefore fundamentally different.
@@
While examining the embedding and attention equations through the lens of basis transformations and perturbative updates, a striking structural parallel emerges between Transformer computation and the ordered operator expansions familiar from many-body physics.
@@
Rather than drawing post-hoc analogies between machine learning and physics, we follow a constructive path where each parallel emerges from the mathematical structure of the architecture itself. Within this formulation, the empirical properties of deep Transformers, including stability at large depth, representational saturation, and multi-head effectiveness—acquire natural structural interpretations. Residual connections generate an exact algebraic expansion over ordered interaction paths, normalization regulates the magnitude of successive updates, and multi-head attention implements a channel factorization of the interaction operator. These elements thus emerge as organizing principles of operator composition rather than isolated heuristics.
We assume familiarity with the basic Transformer components [29], specifically self-attention and residual connections. We analyze their canonical forms here, noting that modern architectural variations preserve this core operator structure. While our analysis focuses on autoregressive models, where causal masking strictly enforces non-Hermitian, directional interactions common in generative tasks, bidirectional variants like BERT [7] permit potentially symmetric attention, yielding operators closer to Hermitian but less suited for sequence generation.
@@ II The Kinematics: Embeddings as Basis Transformation
The Transformer begins from a mathematically analogous construction. A token enters the model as a one-hot vector: a discrete index that labels vocabulary position but carries no geometry, analogous to labeling a lattice site. The embedding matrix $W_{E}$ performs a learned linear transformation on this index space, projecting the discrete label $e_{i}$ into a continuous latent vector $x_{i}^{(0)}$of dimension $d_{\mathrm{model}}$ [15, 7]:
$x_{i}^{(0)}=W_{E}e_{i}.$
@@ III The Dynamics: Perturbation as a Mental Picture for Attention
To understand how the Transformer evolves the state of a sequence [29], we must first define its core computational block. The architecture operates through the repeated application of a modular structure: as shown in Fig. 1 (a), it consists of a stack of $L$ identical layers applied sequentially to (b) a system of $N$ degrees of freedom (token states) within a layer.
@@ III.3 Self-Attention as an Effective Interaction
Furthermore, the Transformer’s residual mechanism strictly lacks the imaginary unit, acting instead as a purely real generator. Consequently, the layer update is non-unitary. Rather than exhibiting oscillatory dynamics, representations are iteratively reshaped, making the forward pass mathematically akin to a dissipative flow rather than conservative time evolution.
@@ III.6 From Theory to Engineering Practice
Collectively, these constraints ensure that the residual update $\|\delta x_{l}\|$ in the identity stream $x_{l+1}=x_{l}+\delta x_{l}$ remains small [10, 5], placing the dynamics in a regime where a perturbative depth expansion is a valid approximation.
@@ IV The Evolution of Depth: Time-Ordered Dyson Series
When expanded, this product generates a sum over all possible ordered permutations of the interaction operators [30]:
@@ IV.1 Layer Norm as Wavefunction Renormalization
While high-order perturbation expansions in quantum systems frequently diverge, the Transformer architecture achieves convergence through an iterative ’re-gauging’ mechanism known as Layer Normalization [1]. Physically, this is similar to a dynamic Wavefunction Renormalization applied at each order of the expansion. Rather than allowing the self-energy contributions to shift the state vector’s norm into unphysical regimes, Layer Norm continuously re-scales the ’dressed’ state. This ensures that the iterative Dyson series (represented by the network depth) remains mapped onto a stable, unit-variance manifold, effectively acting as a non-perturbative regulator that preserves the stability of the information flow, as empirically observed in models like GPT-3 [4], where architectures with up to 96 layers maintain coherent gradients during training without explosion. The exact placement of this renormalization step and its precise form vary across modern architectures, but the structural role as a stability regulator of the operator composition remains unchanged.
@@ V CAUSALITY AND MEASUREMENT: THE SPATIAL ARROW AND UNEMBEDDING
The forward pass serves to prepare the boundary state $|\Psi_{N}\rangle$ for measurement. By the final layer $L$, the token at position $N$ has evolved into a fully ”dressed” state, encoding the lattice history through non-local interactions and local renormalizations. Because of this causal structure, $|\Psi_{N}\rangle$ acts as a sufficient representation for the next-token prediction and no global measurement of the full sequence is required.
@@
However, this mapping is physically distinct. While quantum probabilities emerge from the geometry of Hilbert space via squared wavefunction amplitudes, the Transformer’s distribution arises from statistical optimization over a training corpus. This framework emphasizes the shared algebraic sequence—projection followed by normalization—rather than suggesting a common physical origin.
@@ VI CONCLUSION AND OUTLOOK
The operator-theoretic bridge proposed here, however, rests on a structural parallel rather than an exact identity mapping. While physical operators follow from first principles, such as Hermiticity and unitarity, Transformer weights are empirical outcomes of gradient descent and nonlinear in nature. Despite these differing origins, both disciplines confront the same structural bottleneck: the stabilization of large ordered products of non-commuting, non-Hermitian operators against divergence or rank collapse — an instability independent of whether the constituent operators arise from a Hamiltonian or a learned weight matrix.
@@
The transferability of techniques depends on identifying which operator properties govern this composition. These are not Hermiticity or unitarity, but the non-Hermitian, dense, and potentially defective character of individual layers which is precisely the regime addressed by non-Hermitian random matrix theory [20, 25] and pseudospectral analysis [27]. Transformer weights sit within the domain of applications of these tools.
@@
While the constituent layer operations are nonlinear, the depth expansion in Eq. 7 generates a summation over state vectors within the linear representation space. Consequently, sequence-extrapolation methods that operate directly on the trajectory of states, such as DIIS [18] used in fixed-point solvers, or vector Padé [18] and Borel [3] resummation, may prove directly applicable for extracting stable, converged representations from this discrete series.
Pseudospectral analysis therefore offers a more insightful diagnostic of stability than spectral radius alone when analyzing the non-Hermitian operators governing layer updates. The remaining open question is what structure the training process itself imposes on these operators, and whether that structure renders the resulting dynamics significantly more tractable than worst-case non-Hermitian analysis would suggest.
@@
The transfer of knowledge between these fields is inherently reciprocal. From a physics perspective, the empirical stability of large-scale Transformers is remarkable. They maintain coherent state vectors across hundreds of sequential, non-Hermitian layer operations without diverging or collapsing into a single mode, despite being shaped by optimization across hundreds of billions of scalar parameters. Deep learning achieves this not through physical conservation laws, but through structural interventions like residual connections and layer normalization that act as algorithmic regulators for strongly out-of-equilibrium dynamics.
@@
As discussed earlier, deep learning has the mathematical freedom to arbitrarily rescale state vectors to enforce stability that is prohibited by the conservation laws governing physical quantum systems. However, this divergence in constraints is what makes cross-disciplinary exchange valuable. Translating AI’s empirically successful stabilization mechanisms into operator language may offer physicists new perspectives on controlling complex non-Hermitian dynamics. Conversely, the deep learning community can draw upon the analytical frameworks of many-body physics to better characterize the stability and structure of these architectures.
@@
By identifying operator product control as the shared numerical bottleneck, we move beyond analogy toward systematic translation. The value of this perspective lies not in providing immediate solutions, but in establishing that physicists and machine learning researchers confronting similar mathematical and structural bottlenecks — and can therefore begin working on it together.

### 17 Mar 2026 v2

@@ Abstract
Transformer architectures are typically described in algorithmic and statistical terms, leaving their internal mechanics without a familiar structural language for researchers trained in physical theories. To bridge this gap, we develop a complementary operator-theoretic framework that recasts their mechanics in a language familiar to many-body physics. Beginning from the token as a discrete index without intrinsic geometry, we show that embedding corresponds to a basis transformation into a continuous representation space. Once such a reference basis is established, self-attention naturally assumes the role of a non-Hermitian interaction operator, and network depth implements an ordered composition of these interactions. Within this formulation, several empirical properties of deep Transformers—including stability at large depth, representational saturation, and the effectiveness of multi-head decomposition—find natural structural interpretations as consequences of regulated operator composition. Together, channel factorization and normalization emerge as organizing structural logic rather than isolated architectural choices. This perspective does not rely on post-hoc analogy, but follows a constructive path where each parallel arises from the preceding structural step. By recasting Transformer mechanics in operator language, the framework lowers the conceptual barrier between deep learning and many-body physics through shared mathematical structure, making tools and intuitions from each domain more readily legible to the other.
@@ I Introduction: Two Different Universes
Modern many-body physics and deep learning originate from fundamentally different intellectual starting points. In quantum mechanics, a system is bounded by its Hamiltonian $H$ constructed from explicit dynamical laws, which defines its allowable states. Machine-learning systems, in contrast, are discovered through optimization: their structure emerges from data, with parameters adjusted to minimize a global loss over a statistical corpus whose generative rules are not explicitly defined.
This difference in origin is not superficial. Physical operators inherit constraints, such as Hermiticity, unitarity, and conservation laws, which are non-negotiable consequences of physical laws. On the other hand, learned operators inherit no such constraints as properties like stability are enforced through deliberate architectural choices rather than first principles. The degrees of freedom available to each framework are therefore fundamentally different.
@@
Rather than drawing post-hoc analogies between machine learning and physics, we follow a constructive path where each parallel emerges from the mathematical structure of the architecture itself. Within this formulation, the empirical properties of deep Transformers, including stability at large depth, representational saturation, and multi-head effectiveness, acquire natural structural interpretations. Residual connections generate an exact algebraic expansion over ordered interaction paths, normalization regulates the magnitude of successive updates, and multi-head attention implements a channel factorization of the interaction operator. These elements thus emerge as organizing principles of operator composition rather than isolated heuristics.
We assume familiarity with the basic Transformer components [30], specifically self-attention and residual connections. We analyze their canonical Post-Norm forms here, as this provides the cleanest structural picture for the operator-theoretic framework, noting that modern architectural variations preserve these essential mechanisms despite differences in implementation. When invoking causal masking and the transfer-matrix-like generation mechanism, we draw on autoregressive variants whose directional interactions strictly enforce non-Hermitian operator structure. Bidirectional variants like BERT [7] permit potentially symmetric attention, yielding operators closer to Hermitian, but are not the focus here
@@ II The Kinematics: Embeddings as Basis Transformation
The Transformer begins from a mathematically analogous construction. A token enters the model as a one-hot vector: a column vector of dimension $V$ with zeros everywhere except a single entry of one at the index corresponding to its vocabulary position, carrying no geometry beyond that discrete label, analogous to labeling a lattice site. The embedding matrix $W_{E}$ performs a learned linear transformation on this index space, projecting the discrete label $e_{i}$ into a continuous latent vector $x_{i}^{(0)}$of dimension $d_{\mathrm{model}}$ [15, 7]:
$\underbrace{x_{i}^{(0)}}_{\mathbb{R}^{d_{\text{model}}}}=\underbrace{W_{E}}_{d_{\text{model}}\times V}\underbrace{e_{i}}_{\mathbb{R}^{V}}\textbf{}$ (1)
@@ III The Dynamics: Perturbation as a Mental Picture for Attention
To understand how the Transformer evolves the state of a sequence [30], we must first define its core computational block. The architecture operates through the repeated application of a modular structure: as shown in Fig. 1 (a), it consists of a stack of $L$ structurally identical layers, each with distinct learned weights, applied sequentially to (b) a system of $N$ degrees of freedom (token states) within a layer.
@@ III.3 Self-Attention as an Effective Interaction
Furthermore, the Transformer’s residual mechanism strictly lacks the imaginary unit, acting instead as a purely real generator. Consequently, the layer update is non-unitary. Rather than exhibiting oscillatory dynamics, representations are iteratively reshaped through a non-unitary, irreversible forward pass.
@@ III.6 From Theory to Engineering Practice
Residual connections implement the identity shortcut [10], with normalization further stabilizing the operator composition throughout training [5], collectively ensuring a well-regulated forward pass.
@@ IV The Evolution of Depth: Time-Ordered Dyson Series
When expanded, this product generates a sum over all possible ordered compositions of the interaction operators [31]:
@@
It is interesting to note that state-dependent operators of this kind have a natural counterpart in time-dependent density functional theory (TDDFT) [20], where the effective Hamiltonian $\hat{H}[\rho(t)]$ similarly depends on the evolving state at each time step. The standard Transformer in fact shares both features, state-dependent operators and layer-varying parameters, making TDDFT a potentially richer analogy than the static Dyson series. However, as the goal of this work is to establish a clear structural foundation, we retain the simpler time-ordered operator picture and leave the TDDFT correspondence to future investigation.
@@ IV.1 Layer Norm as Wavefunction Renormalization
While high-order perturbation expansions in quantum systems frequently diverge, The Transformer architecture helps maintain stability through an iterative “re-gauging” mechanism known as Layer Normalization [1]. Physically, this plays a role analogous to a renormalization-like rescaling applied at each order of the expansion. Rather than allowing successive operator compositions to drive the state vector’s norm into uncontrolled regimes, LayerNorm continuously rescales the dressed state, keeping the depth expansion within a bounded region of representation space.
It effectively acts as a stability regulator for the operator composition, empirically observed in by models like GPT-3 [4], where architectures with up to 96 layers maintain coherent representations across depth. The exact placement of this renormalization step and its precise form vary across modern architectures, but the structural role as a stability regulator of the operator composition remains unchanged.
@@ V CAUSALITY AND MEASUREMENT: THE SPATIAL ARROW AND UNEMBEDDING
The forward pass serves to prepare the boundary state $|\Psi_{N}\rangle$ for measurement. By the final layer $L$, the token at position $N$ has evolved into a fully ”dressed” state, encoding the contextual influence of all preceding tokens through non-local interactions and local renormalizations. Because of this causal structure, $|\Psi_{N}\rangle$ acts as a sufficient representation for the next-token prediction and no global measurement of the full sequence is required.
@@
However, this mapping is physically distinct. While quantum probabilities emerge from the geometry of Hilbert space via squared wavefunction amplitudes, This framework emphasizes the shared algebraic sequence, projection followed by normalization, rather than suggesting a common physical origin.
@@ VI CONCLUSION AND OUTLOOK
The operator-theoretic bridge proposed here, however, rests on a structural parallel rather than an exact identity mapping. While physical operators follow from first principles, such as Hermiticity and unitarity, Transformer weights are empirical outcomes of gradient descent and nonlinear in nature. Despite these differing origins, both disciplines confront similar structural bottleneck: the stabilization of large ordered products of non-commuting, non-Hermitian operators against divergence or rank collapse — an instability independent of whether the constituent operators arise from a Hamiltonian or a learned weight matrix.
@@
The transferability of techniques depends on identifying which operator properties govern this composition. These are not Hermiticity or unitarity, but the non-Hermitian, dense, and potentially defective character of individual layers. The matrices underlying these transformations share structural features with the classes of operators studied in non-Hermitian random matrix theory [21, 26] and pseudospectral analysis [28]. While the present work does not establish a direct mapping, these frameworks may provide natural tools for diagnosing stability and sensitivity in deep operator compositions.
@@
While the constituent layer operations are nonlinear, the depth expansion in Eq. 8 generates a summation over state vectors within the linear representation space. For architectures that seek fixed points, such as Deep Equilibrium Models [2], fixed-point acceleration methods such as Anderson mixing, the numerical analog of DIIS [18], are already employed in practice. For standard Transformers with depth-varying operators, sequence extrapolation methods operating directly on the state trajectory, such as vector Padé and Borel [3] resummation, may offer complementary tools for extracting stable representations from this discrete series, though their applicability to nonlinear compositions requires further investigation.
@@
The transfer of knowledge between these fields is inherently reciprocal. From a physics perspective, the empirical stability of large-scale Transformers is remarkable. They maintain coherent state vectors across many sequential, non-Hermitian layer operations without catastrophic divergence in practice, despite being shaped by optimization across billions of scalar parameters. Deep learning achieves this not through physical conservation laws, but through structural interventions like residual connections and layer normalization that act as algorithmic regulators for strongly out-of-equilibrium dynamics.
@@
As discussed earlier, deep learning has the mathematical freedom to rescale state vectors through architectural choices to enforce stability that is prohibited by the conservation laws governing physical quantum systems. However, this divergence in constraints is what makes cross-disciplinary exchange valuable. Translating AI’s empirically successful stabilization mechanisms into operator language may offer physicists new perspectives on controlling complex non-Hermitian dynamics. Conversely, the deep learning community can draw upon the analytical frameworks of many-body physics to better characterize the stability and structure of these architectures.
@@
By identifying operator product control as the shared numerical bottleneck, we move beyond analogy toward systematic translation. The value of this perspective lies not in providing immediate solutions, but in establishing that physicists and machine learning researchers confronting similar mathematical and structural bottlenecks, and can therefore begin working on them together.